# From CVE Entries to Verifiable Exploits: An Automated Multi-Agent Framework for Reproducing CVEs

### v1

@@ 3.2 Builder
Once the knowledge base is populated and the vulnerable version of the project’s source code is downloaded, the next objective is to build the project in a way that allows the exploit to be executed. To achieve this, our pipeline employs two developer LLM agents: the Pre-requisite Developer Agent, which analyzes and plans the environment setup, and the Setup Developer Agent, which executes this plan to configure the vulnerable environment. As well as one critic, the Setup Critic Agent, which analyzes the logs of the Setup Developer and evaluates the setup for the project, allowing the system to correct its mistakes. Below, we discuss the functionality of these agents.
3.2.1 Pre-Requisite Developer Agent 3
Tools. For this step, we only provide the LLM agent with access to read-only tools, i.e., execute_ls_command, and get_file_by_name, which the agent can use to traverse the code base and read files, because we do not want the agent to execute any commands or write to files in this phase.
Goal/Rationale. We introduce this agent as an initial exploration and planning step before actually setting up the vulnerable project. This is done for three key reasons.
(1) Our preliminary experiments reveal that many projects contain inaccurate and incomplete information in their README files (e.g., vulnerable version v1.2.7 of lunary for CVE-2024-5129, includes incorrect paths to .env files that are crucial for the setup). Without the Pre-Requisite agent, these details were consistently overlooked by LLMs, causing setup failures. Identifying and correcting such issues early reduces unsuccessful attempts, conserves the LLM’s context window, and improves the likelihood of successful environment setup.
(2) During environment setup, the agent’s job is to explore and analyze the source code to identify critical components and to build the project by executing a series of commands. For large projects, having to do both tasks with one agent can exceed the context limit and reduce efficiency (as demonstrated in the ablation study in Section 4.5). To address this, we delegate the exploration of code base to a dedicated agent, the Pre-Requisite Developer, allowing the setup process to focus more on building the project.
@@ 3.2.2 Setup Developer Agent 4
Tools. For this step, we allow access to the following tools, i.e., get_file, write_to_file, execute_ls_command, execute_linux_command, and set_environment_variable.
Goal/Rationale. The Setup Developer begins in the directory containing the vulnerable version’s source code and focuses on configuring the project based on instructions from the Pre-Requisite Developer, exploring the codebase when necessary. For CVEs involving third-party libraries, our initial experiments showed that setup is greatly simplified by using package managers like pip or npm to install specific vulnerable versions instead of building from source. For example, in the case of CVE-2024-4340 affecting sqlparse, the Pre-Requisite Developer instructed the installation of version 0.4.4, and the Setup Developer executed pip install sqlparse==0.4.4 directly (command 4a, Figure 2), resulting in a smooth setup. If the package manager fails, the Setup Developer falls back to building from source. Typically, the Pre-Requisite Developer also recommends running a basic PoC to confirm readiness before handing off to the next agent (e.g., commands 4e and 4f in Figure 2).
@@ 3.2.3 Setup Critic Agent 5
Goal/Rationale. The Setup Critic evaluates whether the project setup performed by the Setup Developer is correct and complete, ensuring that the environment is properly configured for the vulnerability described in the CVE knowledge base to be exploited. Additionally, this agent plays a critical role in detecting deceptive or shortcut behaviors by the LLM, such as fabricating a mock project instead of performing a genuine setup, or deliberately injecting vulnerabilities into the codebase to simplify exploitation. These undesirable patterns are discussed in detail in Section 4.1.1, which informs the instruction design for the Setup Critic. Based on this, the critic is equipped to identify both functional and security-related flaws in the setup process.
@@
1
2"""
3PoC CVE-2024-4340 (sqlparse < 0.5.0)
[…]
5python3 exploit.py 10000
6"""
7import sys
8import sqlparse
[…]
10if len(sys.argv) != 2:
11print("Usage: python3 exploit.py <depth>")
12sys.exit(1)
13arg = sys.argv[1]
14try:
15depth = int(arg)
16payload = "[" * depth + "]" * depth
17except ValueError:
18payload = arg
19sqlparse.parse(payload)
[…]
21main()
@@
1
2Traceback (most recent call last):
3File "poc.py", line 20, in <module>
4sqlparse.parse(payload)
5....
6File "sqlparse/sql.py", line 214, in flatten
7yield from token.flatten()
8[Previous line repeated 983 more times]
9RecursionError: maximum recursion depth exceeded
@@ 3.3 Exploiter
After successfully configuring the vulnerable system, the next step in our pipeline is to generate an exploit for the given vulnerability. CVE-Genie accomplishes this using two LLM agents: the Exploit Developer, which develops the PoC for an exploit in the vulnerable environment, and the Exploit Critic, which analyzes the logs of the Exploit Developer and evaluates the exploit for the given vulnerability. Below, we discuss the detailed functionality of these two agents.
@@ 3.3.1 Exploit Developer Agent 6
Goal/Rationale. The Exploit Developer is responsible for crafting and demonstrating exploits for vulnerabilities listed in the CVE knowledge base on pre-configured vulnerable systems. If a PoC is provided, the agent replicates and verifies it by executing the script with the appropriate triggering input before delivering the final, working PoC. For example, in CVE-2024-4340 (Figure 2), the PoC was available, so the agent validated the setup (6b), demonstrated the exploit (6c) by triggering RecursionError in sqlparse/sql.py (see Figure 3(b)), and submitted the verified PoC script (see Figure 3(a)). If no PoC or exploitation steps are available, the Exploit Developer iteratively analyzes the codebase using available tools to develop a functional exploit. Upon success, it produces a Python PoC script that accepts the crashing input via command-line, includes comments on expected input format, and provides an example input. This ensures that the CTF Verifier can reliably validate the exploit without repeating the analysis.
@@ 3.3.2 Exploit Critic Agent 7
Similar to the role of the Setup Critic, the Exploit Critic is tasked with evaluating the behavior of the Exploit Developer by analyzing its execution logs. Its primary goal is to determine whether the generated exploit is both valid and of high quality. This includes verifying that the exploit could plausibly succeed in a real-world setting; that it does not rely on manipulating or altering the original setup in illegitimate ways; and that it avoids fabricated shortcuts, oversimplified logic, or artificial assumptions that would not hold in practice.
To perform this assessment, the Exploit Critic leverages a predefined set of behavioral patterns and common failure cases, as detailed in Section 4.1.2. By comparing the exploit’s behavior against these known issues, the critic can detect actions that are unfaithful to the original intent or technically incorrect. When flaws are found, the Exploit Critic provides structured, actionable feedback designed to help improve the robustness and realism of the exploit in subsequent iterations.
@@ 3.4 CTF Verifier
1
2import os, subprocess, sys
3FLAG = "3xploit66full"
[…]
5# Configuring PYTHONPATH
[…]
7cmd = [sys.executable, "exploit.py", "10000"]
8return subprocess.run(cmd, timeout=20)
[…]
10output = (proc.stdout or "") + (proc.stderr or "")
11return proc.returncode != 0 and "RecursionError" in output
[…]
13pre_setup()
14try: result = run_exploit()
15except subprocess.TimeoutExpired: sys.exit(1)
16if post_exploit(result):
17print(FLAG)
@@
1
2import os, sys, importlib, traceback
3FLAG = "3xploit66full"
[…]
5assert sqlparse.__version__ < "0.5.0"
[…]
7sys.argv = ["exploit.py", "10000"]
8try:
9importlib.import_module("exploit").main()
10return "no_exception", ""
11except RecursionError: return "recursion_error", traceback.format_exc()
12except: return "other_exception", traceback.format_exc()
[…]
14return status == "recursion_error" and "sqlparse/sql.py" in tb
[…]
16pre_setup()
17status, tb = run_exploit_inprocess()
18if exploit_succeeded(status, tb):
19print(FLAG)
@@ 3.4.1 Verifier Developer Agent 8
Goal/Rationale. Following the CTF methodology, we prompt the Verifier Developer to create a general CTF-style Python verifier script for a given PoC script. So, if the verifier script runs the PoC and it successfully triggers the vulnerability, it returns this flag (3xploit66ful). We prompt the Verifier Developer to format the verifier script in three structured steps:
@@
    Post-Setup: Verifies the success of the exploit and if successful, the script returns the flag.
This structured format ensures consistent and reliable validation across a broad spectrum of vulnerability exploit types.
@@ 3.4.2 Flag Checker 9
Once the verifier script is generated for the given exploit, before passing it to the critic agent, the Flag Checker executes the verifier script to ensure that it runs with the exploit script without any errors and produces the expected output i.e., the correct flag. For instance, figure 2 shows the correct flag 3xploit66ful is released when executing the verifier scripts generated for CVE-2024-4340. If a script fails due to a runtime error, or the flag is missing from the output, the system assumes that the verifier is incorrect. Feedback is then sent to the Verifier Developer to revise the script. This process is repeated iteratively, with a maximum of five attempts allowed to achieve functional correctness of the verifier script.
@@ 3.4.3 Verifier Critic Agent 10
Once the verifier and exploit scripts pass functional validation via the Flag Checker, they are further examined by the Verifier Critic, which performs a similar evaluative role as the other critic agents, focusing specifically on the response produced by the Verifier Developer. It examines whether the verification process was properly carried out, without assumptions, omissions, or unreliable heuristics. Using the behavioral issues identified in Section 4.1.3, the agent flags incomplete or misleading verification strategies and provides feedback on how the verification could be made more robust or accurate. For instance, Figure 4 shows a detailed example of verifiers generated for CVE-2024-4340 with a weak and bypassable verification criteria, which was corrected using the Verifier Critic’s feedback.
@@ 3.4.4 Store Reproduced CVE 11
If the given CVE passes the Verifier Critic check, CVE-Genie marks it as reproduced and stores the VM snapshot for the vulnerable environment as well as the exploit and verifier scripts. Moreover, CVE-Genie also stores the metadata such as LLM cost, time spent, and all agents conversations.
4 Experimental Methodology
We begin by identifying the most suitable LLMs for integration into CVE-Genie by evaluating ten leading models (Table 3) across CVE-Genie’s three core modules, Builder, Exploiter, and CTF Verifier, using a diverse set of 60 post-knowledge-cutoff CVEs (Section 4.1). Based on this evaluation, we select the optimal LLM configuration and conduct a four-part evaluation of CVE-Genie: (1) a baseline study assessing consistency and efficiency across various vulnerabilities, languages, and projects (Section 4.2); (2) robustness testing under incomplete CVE information (Section 4.3); (3) large-scale evaluation on 841 CVEs (Section 4.4); and (4) an ablation study to assess the impact of key components of CVE-Genie on its scalability and effectiveness (Section 4.5). These evaluations are guided by the following research questions:
  • RQ1: How does the performance of CVE-Genie vary with the types of vulnerabilities, languages, and projects?
  • RQ2: Can CVE-Genie effectively reproduce CVEs with limited information, and which CVE report components are most crucial for reproduction?
  • RQ3: Can CVE-Genie reproduce a broad range of CVEs across diverse CWEs, languages, and projects?
  • RQ4: Is the complex architecture of CVE-Genie necessary, or can standalone LLMs achieve comparable results?
Model # Params Max Input Max Output Reas-oning Open Src Cutoff
@@ 4.1 Selecting Optimal LLMs for CVE-Genie
Methodology. CVE-Genie is composed of four modules. The Processor’s LLM-based component, the Knowledge Builder, only performs a trivial task of summarization, so we assign a lightweight LLM (o4-mini) to it. In contrast, the remaining modules i.e., Builder, Exploiter, and CTF Verifier, require strong security vulnerability reasoning and SWE capabilities. Thus for each of these modules, we systematically select the best performing LLM for “developer” and “critic” tasks using the following 5-step procedure:
  1. 1.
    Candidate Selection: For each module, we identify subsets of LLMs from Table 3 as potential developers and critics, based on prior evidence of their capabilities, as well as time and cost constraints.
  2. 2.
    Developer Evaluation: Each candidate developer LLM is run for the given module, and its outputs are manually scored by the authors. The highest-scoring LLM is chosen as the developer.
  3. 3.
    Critic Prompt Design: We analyze common mistakes made by developers and use these insights to design a task-specific critic prompt for the module, enabling the critic to effectively detect those mistakes.
  4. 4.
    Critic Evaluation: Each candidate critic LLM is evaluated on recall or true positive rate (TPR) and specificity or true negative rate (TNR). When TPR is equal, we prefer the model with higher TNR, and vice versa. The critic with the best trade-off is selected.
  5. 5.
    Feedback Loop Assessment: We rerun the chosen best developer–critic pair for the given module, and evaluate the correctness of critic’s feedback and the developer’s capability to correctly address the critic’s feedback.

### v2

@@ 3.2 Builder
Once the knowledge base is populated and the vulnerable version of the project’s source code is downloaded, the next objective is to build the project in a way that allows the exploit to be executed. To achieve this, our pipeline employs two developer LLM agents: the Pre-Requisite Developer Agent, which analyzes and plans the environment setup, and the Setup Developer Agent, which executes this plan to configure the vulnerable environment. As well as one critic, the Setup Critic Agent, which analyzes the logs of the Setup Developer and evaluates the setup for the project, allowing the system to correct its mistakes. Below, we discuss these in detail.
3.2.1 Prerequisite Developer Agent 3
We introduce this agent as an initial exploration and planning step before actually setting up the vulnerable project. This is done for three key reasons.
(1) Our preliminary experiments reveal that many projects contain inaccurate and incomplete information in their README files (e.g., vulnerable version v1.2.7 of lunary for CVE-2024-5129, includes incorrect paths to .env files that are crucial for the setup). Without the Prerequisite agent, these details were consistently overlooked by LLMs, causing setup failures. Identifying and correcting such issues early reduces unsuccessful attempts, conserves the LLM’s context window, and improves the likelihood of successful setup.
(2) During environment setup, the agent’s job is to explore and analyze the source code to identify critical components and to build the project by executing a series of commands. For large projects, having to do both tasks with one agent can exceed the context limit and reduce efficiency (as demonstrated in the ablation study in Section 4.2). To address this, we delegate the exploration of the code base to a dedicated agent, the Pre-Requisite Developer, allowing the setup process to focus more on building the project.
@@ 3.2.1 Prerequisite Developer Agent 3
Tools. For this step, we only provide the LLM agent with access to read-only tools, i.e., execute_ls_command, and get_file_by_name, which the agent can use to traverse the code base and read files, because we do not want the agent to execute any commands or write to files in this phase (see Appendix E for tools details).
@@ 3.2.2 Setup Developer Agent 4
The Setup Developer begins in the directory containing the vulnerable version’s source code and focuses on configuring the project based on instructions from the Pre-Requisite Developer, exploring the codebase when necessary. For CVEs involving third-party libraries, our initial experiments showed that setup is greatly simplified by using package managers like pip or npm to install specific vulnerable versions instead of building from source. For example, in the case of CVE-2024-4340 affecting sqlparse, the Pre-Requisite Developer instructed the installation of version 0.4.4, and the Setup Developer executed pip install sqlparse==0.4.4 directly (command 4a, Figure 2), resulting in a smooth setup. If the package manager fails, the Setup Developer falls back to building from source. Typically, the Pre-Requisite Developer also recommends running a basic PoC to confirm readiness before handing off to the next agent (e.g., commands 4e and 4f in Figure 2).
Tools. For this step, we allow access to the following tools, i.e., execute_ls_command, execute_linux_command, get_file, write_to_file, and set_environment_variable.
@@ 3.2.3 Setup Critic Agent 5
The Setup Critic evaluates whether the project setup performed by the Setup Developer is correct and complete, ensuring that the environment is properly configured for the vulnerability described in the CVE knowledge base to be exploited. Additionally, this agent plays a critical role in detecting deceptive or shortcut behaviors by the LLM, such as fabricating a mock project instead of performing a genuine setup, or deliberately injecting vulnerabilities into the codebase to simplify exploitation. These undesirable patterns are discussed in detail in Section 3.5.1, which informs the instruction design for the Setup Critic. Based on this, the critic is equipped to identify both functional and security-related flaws in the setup process.
@@
1"""
2PoC - CVE-2024-4340 (sqlparse < 0.5.0)
3EXAMPLE THAT CRASHES v0.4.4
4    python3 exploit.py 10000
5"""
6import sys
7import sqlparse
8def main() -> None:
9 if len(sys.argv) != 2:
10 print("Usage: python3 exploit.py <depth>")
11 sys.exit(1)
12 arg = sys.argv[1]
13 try:
14 depth = int(arg)
15        payload = "[" * depth + "]" * depth
16 except ValueError:
17 payload = arg
18    sqlparse.parse(payload)
19if __name__ == "__main__":
20 main()
@@
1Traceback (most recent call last):
2 File "poc.py", line 20, in <module>
3 sqlparse.parse(payload)
4 ....
5 "sqlparse/sql.py", line 214, in flatten
6 yield from token.flatten()
7 [Previous~line~repeated~983~more~times]
8RecursionError: maximum recursion depth exceeded
@@ 3.3 Exploiter
After successfully configuring the vulnerable system, the next step in our pipeline is to generate an exploit for the given vulnerability. CVE-Genie accomplishes this using two LLM agents: the Exploit Developer, which develops the PoC for an exploit in the vulnerable environment, and the Exploit Critic, which analyzes the logs of the Exploit Developer and evaluates the exploit for the given vulnerability. Below, we discuss their detailed functionalities.
@@ 3.3.1 Exploit Developer Agent 6
The Exploit Developer is responsible for crafting and demonstrating exploits for vulnerabilities listed in the CVE knowledge base on pre-configured vulnerable systems. If a PoC is provided, the agent replicates and verifies it by executing the script with the appropriate triggering input before delivering the final, working PoC. For example, in CVE-2024-4340 (Figure 2), the PoC was available, so the agent validated the setup (6b), demonstrated the exploit (6c) by triggering RecursionError in sqlparse/sql.py (see Figure 3(b)), and submitted the verified PoC script (see Figure 3(a)). If no PoC or exploitation steps are available, the Exploit Developer iteratively analyzes the codebase using available tools to develop a functional exploit. Upon success, it produces a Python PoC script that accepts the crashing input via the command line, includes comments on the expected input format, and provides an example input. This ensures that the CTF Verifier can reliably validate the exploit without repeating the analysis.
@@ 3.3.2 Exploit Critic Agent 7
Similar to the role of the Setup Critic, the Exploit Critic is tasked with evaluating the behavior of the Exploit Developer by analyzing its execution logs. Its primary goal is to determine whether the generated exploit is both valid and high-quality. This includes verifying that the exploit could plausibly succeed in a real-world setting; that it does not rely on manipulating or altering the original setup in illegitimate ways; and that it avoids fabricated shortcuts, oversimplified logic, or artificial assumptions that would not hold in practice.
To perform this assessment, the Exploit Critic leverages a predefined set of behavioral patterns and common failure cases, as detailed in Section 3.5.2. By comparing the exploit’s behavior against these known issues, the critic can detect actions that are unfaithful to the original intent or technically incorrect. When flaws are identified, the Exploit Critic provides structured, actionable feedback to improve the robustness of the exploit in subsequent iterations.
@@ 3.4 CTF Verifier
1import os, subprocess, sys
2FLAG = "3xploit66full"
3
4def pre_setup():
5    # Configuring PYTHONPATH
6
7def run_exploit():
8 cmd = [sys.executable, "exploit.py", "10000"]
9    return subprocess.run(cmd, timeout=20)
10
11def post_exploit(proc):
12 output = (proc.stdout or "") + (proc.stderr or "")
13    return proc.returncode != 0 and
14            "RecursionError" in output
15
16if __name__ == "__main__":
17 pre_setup()
18 try: result = run_exploit()
19    except subprocess.TimeoutExpired: sys.exit(1)
20 if post_exploit(result):
21 print(FLAG)
@@
1import os, sys, importlib, traceback
2FLAG = "3xploit66full"
3
4def pre_setup():
5    assert sqlparse.__version__ < "0.5.0"
6
7def run_exploit_inprocess():
8 sys.argv = ["exploit.py", "10000"]
9 try:
10        importlib.import_module("exploit").main()
11 return "no_exception", ""
12    except RecursionError:
13        return "recursion_error", traceback.format_exc()
14 except:
15 return "other_exception", traceback.format_exc()
16
17def exploit_succeeded(status, tb):
18    return status == "recursion_error" and
19            "sqlparse/sql.py" in tb
20
21if __name__ == "__main__":
22 pre_setup()
23 status, tb = run_exploit_inprocess()
24 if exploit_succeeded(status, tb):
25 print(FLAG)
@@ 3.4.1 Verifier Developer Agent 8
Following the CTF methodology, we prompt the Verifier Developer to create a general CTF-style Python verifier script for a given PoC script. So, if the verifier script runs the PoC and it successfully triggers the vulnerability, it returns this flag (3xploit66ful). We prompt the Verifier Developer to format the verifier script in three structured steps:
@@
    Post-Setup: Verifies the success of the exploit, and if successful, the script returns the flag.
This verifier architecture provides consistent and reliable exploit validation across a broad spectrum of vulnerability classes. It supports vulnerabilities with directly observable failure signals, such as a RecursionError in CVE-2024-4340 or an AddressSanitizer: heap-buffer-overflow in CVE-2024-10525, as well as higher-level logic flaws. For example, in the case of the missing-authorization flaw in CVE-2024-5129 (Appendix H), the verifier first inserts a legitimate dataset with a known UUID, then executes the exploit by issuing an unauthenticated deletion request targeting that UUID, and finally confirms successful exploitation by verifying that the dataset has been removed from the database.
@@ 3.4.2 Flag Checker 9
Once the verifier script is generated for the given exploit, before passing it to the critic agent, the Flag Checker executes the verifier script to ensure that it runs with the exploit script without any errors and produces the expected output, i.e., the correct flag. For instance, figure 2 shows that the correct flag 3xploit66ful is released when executing the verifier scripts generated for CVE-2024-4340. If a script fails due to a runtime error or if the flag is missing from the output, the system assumes that the verifier is incorrect. Feedback is then sent to the Verifier Developer to revise the script. This process is repeated until a maximum of five attempts is reached, with each attempt ensuring functional correctness of the verifier.
@@ 3.4.3 Verifier Critic Agent 10
Once the verifier and exploit scripts pass functional validation via the Flag Checker, they are further examined by the Verifier Critic, which performs a similar evaluative role as the other critic agents, focusing specifically on the response produced by the Verifier Developer. It examines whether the verification process was properly carried out, without assumptions, omissions, or unreliable heuristics. Using the behavioral issues identified in Section 3.5.3, the agent flags incomplete or misleading verification strategies and provides feedback on how the verification could be made more robust or accurate. For instance, Figure 4 shows a detailed example of verifiers generated for CVE-2024-4340 with weak and bypassable verification criteria, which were corrected using the Verifier Critic’s feedback.
@@ 3.4.4 Store Reproduced CVE 11
If the given CVE passes the Verifier Critic check, CVE-Genie marks it as reproduced and stores the VM snapshot for the vulnerable environment as well as the exploit and verifier scripts. Moreover, CVE-Genie also stores the metadata, such as LLM cost, time spent, and all agents’ conversations.
3.5 Selecting LLMs and Engineering Prompts for CVE-Genie’s Optimal Performance
Upon finalizing the design of CVE-Genie’s architecture, we perform a systematic LLM model and prompt selection process to identify the best performing LLMs and prompt integrations. We evaluate ten state-of-the-art LLMs (Table 3) across CVE-Genie’s three core modules, Builder, Exploiter, and CTF Verifier, using a diverse set of 60 post–knowledge-cutoff CVEs, while simultaneously engineering and refining task-specific prompts for each module.
Model # Params Max Inp Max Out Reas-oning Open Src Cutoff
@@ 3.5 Selecting LLMs and Engineering Prompts for CVE-Genie’s Optimal Performance
Methodology. CVE-Genie is composed of four modules. The Processor’s LLM-based component, the Knowledge Builder, only performs a trivial task of summarization, so we assign a lightweight LLM (o4-mini) to it. In contrast, the remaining modules, i.e., Builder, Exploiter, and CTF Verifier, require strong security vulnerability reasoning and SWE capabilities. Thus, for each of these modules, we systematically select the best performing LLM for “developer” and “critic” tasks using the following 5-step procedure:
(1) Candidate Selection: For each module, we identify subsets of LLMs from Table 3 as potential developers and critics, based on prior evidence of their capabilities, as well as time and cost constraints.
(2) Developer Evaluation: Each candidate developer LLM is run for the given module, and its outputs are manually scored by the authors. The highest-scoring LLM is chosen as the developer.
(3) Critic Prompt Design: We analyze common mistakes made by developers and use these insights to design a module-specific critic prompt, enabling the critic to effectively detect those mistakes.
(4) Critic Evaluation: Each candidate critic LLM is evaluated on true positive rate (TPR) and true negative rate (TNR). When TPRs are equal, we prefer the model with the higher TNR, and vice versa. Then then critic with the best trade-off is selected.
(5) Feedback Loop Assessment: We rerun the chosen best developer–critic pair for the given module, and evaluate the correctness of the critic’s feedback and the developer’s capability to correctly address the critic’s feedback.